I dont know where to start with this question.

1. calculate the frequency of a light that has the wavelength of 41.5nm

2. what type of radiation is this?
a. violet
b. infrared
c. green
d. ultraviolet
e. blue
Answer #1

1)

Given:

lambda = 41.5 nm = 4.15*10^-8 m

use:

f = c/lambda

=(3.0*10^8 m/s)/(4.15*10^-8 m)

= 7.229*10^15 Hz

Answer: 7.23*10^15 Hz

2)

violet : about 400 nm

indigo : about 445 nm

blue : about 475 nm

green : about 510 nm

yellow : about 570 nm

orange : about 590 nm

red : about 650 nm

41.5 nm is less than wavelength of violet.

So, this is ultraviolet

Answer: d